# Mean-field limits for Consensus-Based Optimization and Sampling

### v2

@@ Abstract
For algorithms based on interacting particle systems that admit a mean-field description, convergence analysis is often more accessible at the mean-field level. In order to transpose convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics. In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling. Our approach requires a generalization of Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems. In addition, we present new results on the well-posedness of the particle systems and their mean-field limit, and provide novel stability estimates for the weighted mean and the weighted covariance.
@@ 1 Introduction
This work focuses on the rigorous derivation of quantitative mean-field limits for recently proposed algorithms known as Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants. The former, initially proposed in [CBO], is an algorithm for global optimization tasks, while the latter, proposed in [CBS-Carrillo2021], can be employed for both sampling and optimization purposes. Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels. This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
@@ 1.1 Consensus-Based Optimization (CBO)
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [CBO-PSO-zero-inertia-1]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical]. The distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large exponential parameter $\beta$. The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
@@
A number of extensions of CBO have been proposed in the literature. As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [CBO-high-dim]. Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [Ha2019, Ha2020ConvergenceAE, HA_Discrete_CBO_hetero_noise]. (On this subject, we point out that the results in [Ha2020ConvergenceAE] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [jin2018random], memory effects [MR4160256], and Lévy noise [MR4553241], as well as extensions to multi-objective optimization [klamroth2022consensusbased, borghi2022adaptive] and global optimization with constraints [CBO-hypersurface-Fornasier_2020, MR4329816, CTV21, MR4538901, MR4404221, MR4337731]. Moreover, a variant tailored for objective functions with several minima was proposed in [bungert], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [CBO-PSO-zero-inertia-1, CBO-PSO-zero-inertia-2]. Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [CBO-survey-Totzeck2022]. Furthermore, a free software implementation of CBO and some of its extensions was recently published [Bailo2024].
@@ 1.2 Consensus-Based Sampling (CBS)
More precisely, the authors of [CBS-Carrillo2021] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$. In addition, it was noticed in [CBS-Carrillo2021] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [CBO]. Since the convergence analysis in [CBS-Carrillo2021] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$. This has so far been left as an open problem. Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [bungert], and for rare event estimation in [2023arXiv230409077A].
@@ 1.3 Propagation of chaos
$\mathrm{d}X^{(j)}_{t}=b\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}t+\sigma\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}W^{(j)},\qquad j\in\llbracket 1,J\rrbracket.$ (1.6)
@@
  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [MR4553241]. In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1. There are, however, several differences between our results and those presented in [MR4553241]. Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [MR3696003, MR2597592] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3). Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [MR4553241].
@@ 1.4 Our contributions
  • We obtain quantitative results on the mean-field limit for CBO and CBS. More precisely, we prove results with an explicit rate of convergence with respect to the number $J$ of particles. Investigating the roles of dimension and other problem parameters on the value of the constant prefactor is left for future work.
@@ 2.2 Well-posedness for CBO and CBS
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.3) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty\qquad\text{and}\qquad\sup_{t\in[0,T]}\Bigl{\lvert}\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{\rvert}<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.4)
and the function $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$. Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C\left([0,T],\mathcal{P}_{p}(\mathbf{R}^{d})\right)$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
$\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}\Bigr{)}+\frac{1}{2}\nabla\cdot\nabla\cdot\Bigl{(}D(\overline{\rho},x)\overline{\rho}\Bigr{)}.$ (2.5)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.5) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,\qquad\sup_{t\in[0,T]}\left\lvert\mathcal{M}_{\beta}(\overline{\rho}_{t})\right\rvert<\infty,\qquad\sup_{t\in[0,T]}\left\lVert\mathcal{C}_{\beta}(\overline{\rho}_{t})\right\rVert<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.6)
and the functions $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$, $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$. Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C([0,T],\mathcal{P}_{p}(\mathbf{R}^{d}))$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
@@ 2.3 Key preparatory lemma
$\mathbf{E}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}$ $\leqslant J^{-r}C_{\rm MZ}(r)\mathbf{E}\left[\left(\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{2}\right)^{\frac{r}{2}}\right]$
$\leqslant J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{r}\right]=J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\Bigl{\lvert}Z_{1}-\mathbf{E}[Z_{1}]\Bigr{\rvert}^{r}\right],$
@@ 2.4 Quantitative mean-field limit for CBO
$\forall j\in\llbracket 1,J\rrbracket,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
@@
$\forall N\in\mathbf{N}^{+},\forall j\in\llbracket 1,J\rrbracket,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
@@
We prove below that the following sharper statement holds. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$, and consider the systems (2.8) and (2.9) with the coefficients given by (2.10). If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
@@
$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\frac{1}{2}}.$
@@
As mentioned in Remark 2.7, we prove the sharp result (2.12). It is sufficient to prove this statement for all $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ (by the assumption on $q$, this interval is nonempty). Indeed, if the statement holds true in this case, then by Jensen’s inequality, it holds for all $\mathfrak{p}\in(0,2\vee p_{\mathcal{M}}]$ that
$\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{\mathfrak{p}}\right]\right)^{\frac{1}{\mathfrak{p}}}$ $\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2\vee p_{\mathcal{M}}}\right]\right)^{\frac{1}{2\vee p_{\mathcal{M}}}}$
$\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}=CJ^{-\min\left\{\frac{1}{2},\frac{q-\mathfrak{p}}{2\mathfrak{p}^{2}},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}},$
@@
$\bigl{\lvert}X_{t}^{(j)}-\overline{X}_{t}^{(j)}\bigr{\rvert}^{p}\leqslant$  $\,2^{p-1}\left\lvert\int_{0}^{t}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
$+2^{p-1}\left\lvert\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$

### v3

@@ Abstract
For algorithms based on interacting particle systems that admit a mean-field description, convergence analysis is often more accessible at the mean-field level. In order to transfer convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics. In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling. Our approach requires a generalization of Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems. In addition, we present new results on the well-posedness of the particle systems and their mean-field limit, and provide novel stability estimates for the weighted mean and the weighted covariance.
@@ 1 Introduction
Particle methods provide a powerful paradigm for solving complex global optimization problems leading to highly parallelizable algorithms. Despite widespread and growing adoption, theory underpinning their behavior has been mainly based on meta-heuristics. In application settings involving black-box procedures, or where gradients are too costly to obtain, one relies on derivative-free approaches instead.
This work focuses on the rigorous derivation of quantitative mean-field limits for the recently proposed derivative-free algorithms Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants. The former, initially proposed in [CBO], is an algorithm for global optimization tasks, while the latter, proposed in [CBS-Carrillo2021], can be employed for both sampling and optimization purposes. Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels, which makes it challenging to rigorously prove propagation of chaos.
CBO and CBS methods are a novel family of algorithms providing an alternative to the popular particle swarm optimization (PSO) algorithm [Kennedy_Eberhart, Kennedy1997ThePS] that is found in many optimization toolboxes. Recent studies have shown that CBO is competitive with particle swarm optimization [totzeck2018numerical] for a number of benchmark functions, and able to solve high-dimensional optimization problems arising from machine learning [CBO-high-dim, MR4329816]. By now, these methods have also been adapted for various other tasks and application settings, including finance [MR4337731], rare event simulation [2023arXiv230409077A] and federated learning [CBO-federal-carrillo2023], to mention just a few. Compared to PSO however, this family of algorithms is amenable to theoretical analysis. Since current theoretical guarantees are only available on the level of the mean-field description, providing quantitative estimates quantifying the distance between the evolution of the interacting particles and the corresponding mean-field probability distribution is essential.
This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
@@ 1.1 Consensus-Based Optimization (CBO)
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [CBO-PSO-zero-inertia-1]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical]. While the consensus point does not necessarily coincide with the global minimizer of the cost function $f$, [carrillo2018analytical, Theorem 4.2] shows that the distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large parameter $\beta$. The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
@@
A number of extensions of CBO have been proposed in the literature. As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [CBO-high-dim]. Using componentwise geometric Brownian motion in this way allows for proving consensus formation directly at the particle level, which was achieved in [Ha2019, Ha2020ConvergenceAE, HA_Discrete_CBO_hetero_noise]. (On this subject, we point out that the results in [Ha2020ConvergenceAE] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [jin2018random], memory effects [MR4160256], and Lévy noise [MR4553241], as well as extensions to multi-objective optimization [klamroth2022consensusbased, borghi2022adaptive] and global optimization with constraints [MR4538901, carrillo2024interacting, CTV21, CBO-hypersurface-Fornasier_2020, beddrich2024constrained, MR4329816, CBO_sphere2, MR4337731]. Moreover, a variant tailored for objective functions with several minima was proposed in [bungert], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [CBO-PSO-zero-inertia-1, CBO-PSO-zero-inertia-2]. Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [CBO-survey-Totzeck2022]. Furthermore, a free software implementation of CBO and some of its extensions was recently published [Bailo2024].
@@ 1.2 Consensus-Based Sampling (CBS)
More precisely, the authors of [CBS-Carrillo2021] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$. In addition, it was noticed in [CBS-Carrillo2021] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [CBO]. Since the convergence analysis in [CBS-Carrillo2021] is confined to the mean-field level, it is of interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$. This has so far been left as an open problem. Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [bungert], and for rare event estimation in [2023arXiv230409077A]. It has also been shown that CBS, together with other sampling methods based on interacting particle systems, can be metropolized to correct the sampling error [sprungk2023metropolisadjustedinteractingparticlesampling].
@@ 1.3 Propagation of chaos
$\mathrm{d}X^{(j)}_{t}=b\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}t+\sigma\left(X^{(j)}_{t},\mu^{J}_{t}\right)\,\mathrm{d}W^{(j)},\qquad j\in\llbracket 1,J\rrbracket,$ (1.6)
@@
  • Shortly before completing this work, we found out that propagation of chaos for a variant of CBO had been proved using a coupling approach very recently in [MR4553241]. In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1, which was developed independently. There are, however, several differences between our results and those presented in [MR4553241]. Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [MR3696003, MR2597592] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a concentration inequality to bound the probability that empirical measures exit a compact set (Subsection 2.3). Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [MR4553241].
  • Finally, we mention the recent preprint [koß2024meanfieldlimitconsensus], which appeared a few months after this one and proves propagation of chaos for consensus-based sampling using a compactness argument, similar to the approach used in [CBO-mfl-Huang2021].
@@ 1.4 Our contributions
  • We obtain quantitative results on the mean-field limit for CBO and CBS. More precisely, we prove results with an explicit rate of convergence with respect to the number $J$ of particles. Investigating the roles of the dimension and other problem parameters on the value of the constant prefactor is left for future work.
@@ 2.2 Well-posedness for CBO and CBS
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.3) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$. Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.3) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$, and it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.4)
Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C\left([0,T],\mathcal{P}_{p}(\mathbf{R}^{d})\right)$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
$\partial_{t}\overline{\rho}=\nabla\cdot\Bigl{(}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}\overline{\rho}\Bigr{)}+\frac{1}{2}\nabla\cdot\nabla\cdot\Bigl{(}D(\overline{\rho},x)\overline{\rho}\Bigr{)},$ (2.5)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.5) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.5) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, and it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.6)
Finally, the function $t\mapsto\overline{\rho}_{t}$ belongs to $C([0,T],\mathcal{P}_{p}(\mathbf{R}^{d}))$ and satisfies the following non-local Fokker–Planck equation in the weak sense:
@@ 2.3 Key preparatory lemma
$\mathbf{E}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}$ $\leqslant J^{-r}C_{\rm MZ}(r)\mathbf{E}\left[\left(\sum_{j=1}^{J}\Bigl{\lvert}Z_{j}-\mathbf{E}[Z_{j}]\Bigr{\rvert}^{2}\right)^{\frac{r}{2}}\right]$
$\leqslant J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}Z_{j}-\mathbf{E}[Z_{j}]\Bigr{\rvert}^{r}\right]=J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\Bigl{\lvert}Z_{1}-\mathbf{E}[Z_{1}]\Bigr{\rvert}^{r}\right],$
@@ 2.4 Quantitative mean-field limit for CBO
$\forall j\in\llbracket 1,J\rrbracket,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}\,\mathrm{d}s+\int_{0}^{t}\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}\mathrm{d}W^{(j)}_{s},$ (2.8)
@@
$\forall N\in\mathbf{N}^{+},\forall j\in\llbracket 1,J\rrbracket,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}s+\int_{0}^{t}\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
@@
We prove below the following more general statement: suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$. If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
@@
$W_{p}\Bigl{(}\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\Bigr{)}\leqslant CJ^{-\frac{1}{2}},$
@@
As mentioned in Remark 2.7, we prove the general result (2.12). It is sufficient to prove this statement for all $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ (by the assumption on $q$, this interval is nonempty). Indeed, if the statement holds true in this case, then by Jensen’s inequality, it holds for all $\bar{p}\in(0,2\vee p_{\mathcal{M}}]$ that
$\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{\bar{p}}\right]\right)^{\frac{1}{\bar{p}}}$ $\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2\vee p_{\mathcal{M}}}\right]\right)^{\frac{1}{2\vee p_{\mathcal{M}}}}$
$\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}=CJ^{-\min\left\{\frac{1}{2},\frac{q-\bar{p}}{2\bar{p}^{2}},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}},$
@@
$\bigl{\lvert}X_{t}^{(j)}-\overline{X}_{t}^{(j)}\bigr{\rvert}^{p}\leqslant$  $\,2^{p-1}\left\lvert\int_{0}^{t}b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}s\right\rvert^{p}$
$+2^{p-1}\left\lvert\int_{0}^{t}\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W_{s}\right\rvert^{p}.$